OpenStudy (anonymous):

Please show all work The formula to find the surface area of a sphere is to take the radius of the sphere and square it and multiply that amount by pi and then multiply that by 4. This is written algebraically as S=4πr2, where S is the surface area, r is the radius, and the value of π is 3.14. Rewrite the equation to solve for the radius of the sphere if you know the sphere’s surface area. Then estimate the radius of a sphere given a surface area of 500 meters2.

OpenStudy (zzr0ck3r):

\(S=4\pi r^2\) divide everything by \(4\pi\) \(\frac{S}{4\pi}=r^2\) take the positive square root of both sides, (we are working with length so negative does not make sense) so \(\sqrt{\frac{S}{4\pi}}=r\) now let \(S=500\)

OpenStudy (anonymous):

thank you for the answer

OpenStudy (isaiah.feynman):

I hope you did understand it? @bigpapa021201

OpenStudy (anonymous):

yes I did